evaluate the following integral
∫ a / b + ce^x dx

Grade:12

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:
Hello student,
Please find the answer to your question below

I = \int \frac{a}{b+ce^{x}}dx
I = \int \frac{ae^{-x}}{be^{-x}+c}dx
be^{-x}+c = t
-be^{-x}dx = dt
I = -\frac{a}{b}\int \frac{1}{t}dt
I = -\frac{a}{b}log(t) + d
I = -\frac{a}{b}log(be^{-x}+c) + d